4. Given are the following equations for lentils:


Q_{s} = 2P


and Qd = 300 - P


a. Government imposes a price ceiling of Rs.50 on lentils. Calculate the consumer and producer surplus at the ceiled price?


b. Calculate the dead weight loss.




1
Expert's answer
2022-01-17T12:27:11-0500

Qs = 2P or Ps = 0.5Q, Qd = 300 - P or Pd = 300 - Q.

In equilibrium Qd = Qs, so:

300 - P = 2P,

P = 100,

Q = 2×100 = 200 units.


a. If government imposes a price ceiling of Rs.50 on lentils, then the consumer and producer surplus at the ceiled price are:

Q = 2×50 = 100 units.

"CS = 100\u00d7((300 - 50) + (250 - 50))\/2 = 22,500."

"PS = 0.5\u00d7100\u00d750 = 2,500."


b. The dead weight loss is:

"DWL = 0.5\u00d7(200 - 100)\u00d7(250 - 50) = 10,000."